# Prompt Decomposition and Reconstruction Makes Powerful LLM Jailbreakers

### v2

@@ Abstract
The safety alignment of Large Language Models (LLMs) is vulnerable to both manual and automated jailbreak attacks, which adversarially trigger LLMs to output harmful content. However, current methods for jailbreaking LLMs, which nest entire harmful prompts, are not effective at concealing malicious intent and can be easily identified and rejected by well-aligned LLMs. This paper discovers that decomposing a malicious prompt into separated sub-prompts can effectively obscure its underlying malicious intent by presenting it in a fragmented, less detectable form, thereby addressing these limitations. We introduce an automatic prompt Decomposition and Reconstruction framework for jailbreak Attack (DrAttack). DrAttack includes three key components: (a) ‘Decomposition’ of the original prompt into sub-prompts, (b) ‘Reconstruction’ of these sub-prompts implicitly by in-context learning with semantically similar but harmless reassembling demo, and (c) a ‘Synonym Search’ of sub-prompts, aiming to find sub-prompts’ synonyms that maintain the original intent while jailbreaking LLMs. An extensive empirical study across multiple open-source and closed-source LLMs demonstrates that, with a significantly reduced number of queries, DrAttack obtains a substantial gain of success rate over prior SOTA prompt-only attackers. Notably, the success rate of 78.0% on GPT-4 with merely 15 queries surpassed previous art by 33.1%. Project is available at https://github.com/xirui-li/DrAttack.
Machine Learning, Large Language Model, Jailbreak, Prompt Decomposition, Adversarial Robustness, Defense, Attack, Trustworthy ML, ML Safety
@@ 1 Introduction
The development of large language models (LLMs) (Floridi & Chiriatti, 2020; Touvron et al., 2023; Chowdhery et al., 2023) has enabled AI systems to achieve great success in various tasks. However, these successes are accompanied by emerging vulnerabilities in LLMs, such as susceptibility to jailbreaking attacks (Mozes et al., 2023), which have been rigorously studied (Ouyang et al., 2022; Bai et al., 2022; Glaese et al., 2022; Wolf et al., 2023). To hide harmful intentions of an attack prompt, current automatic jailbreak attacks focus on generating surrounding contents, including suffix contents optimization (Zou et al., 2023; Zhu et al., 2023; Liu et al., 2023; Lapid et al., 2023; Shah et al., 2023), Prefix/system contents optimization (Liu et al., 2023; Huang et al., 2023) and hybrid contents (Li et al., 2023; Ding et al., 2023; Chao et al., 2023; Wei et al., 2023b). However, using a malicious prompt as a single entity makes the inherent malice detectable, undermining the attack success rate.
In this study, we explore a more effective automatic jailbreak strategy that modifies sub-prompts of the original attack prompt to conceal the malice and, at the same time, shrink the search space to make the search more efficient. This strategy is motivated by the insight that attack prompts’ malicious intentions are built on phrases’ semantic binding, e.g., “make a bomb”. In contrast, a separate phrase such as “make” or “a bomb” is more neutral and is less likely to trigger the rejection by LLMs.
To achieve this process, as illustrated in Figure 1, our algorithm, Decomposition-and-Reconstruction Attack (DrAttack), operates through a three-step process: (1) Decomposition via Semantic-Parsing breaks down a malicious prompt into seemingly neutral sub-prompts. (2) Implicit Reconstruction via In-Context Learning reassembles sub-prompts by benign and semantic-similar assembling context. (3) Sub-prompt Synonym Search shrinks the search space to make the search more efficient than optimization in whole vocabulary. Our extensive empirical experiments demonstrate that DrAttack can be effectively applied to a broad class of open-source and closed-source LLMs and substantially increase the success rate over prior SOTA attacks. On GPT-4, DrAttack achieves an attack success rate of over 84.6% (under human evaluation) and 62.0% (under LLM evaluation), surpassing the previous best results by over 30% in absolute terms for both assessments.
Figure 1: An illustration of DrAttack. Attacks by a malicious prompt on LLMs would be rejected (blue). However, with DrAttack’s prompt decomposition and reconstruction with ICL given benign demo (green), the resulting prompt can circumvent LLM’s security measures and generate a harmful response (red). Colored words are sub-prompts.
@@ 2 Related Work
Effective attack techniques that circumvent LLM’s safety detectors uncover the vulnerabilities of LLMs, which could be regarded as a critical process in enhancing the design of safer systems. This is achieved by generating surrounding content to hide the harmful intention of the original prompt. Existing attackers can be roughly categorized into three groups: 1). Suffix-based methods augment the harmful prompt with a suffix content, optimized to trick LLM into generating desired malicious responses (Zou et al., 2023; Zhu et al., 2023; Shah et al., 2023; Lapid et al., 2023). Following GCG (Zou et al., 2023) that optimizes suffixes from random initialization, AutoDAN-b (Zhu et al., 2023) further improves the interpretability of generated suffixes via perplexity regularization. To reduce the dependency on white-box models of these methods, several attempts are made to extend them to black box settings either via words replacement (Lapid et al., 2023) or using open-source LLMs as the surrogate approximator (Shah et al., 2023) 2). Prefix-based methods prepend a "system prompts" instead to bypass the build-in safety mechanism (Liu et al., 2023; Huang et al., 2023). For instance, AutoDAN-a (Liu et al., 2023) searches for the optimal system prompt using a genetic algorithm. 3). Hybrid methods insert several tokens to let the harmful prompts be surrounded with benign contexts ("scenarios")  (Yu et al., 2023; Li et al., 2023; Ding et al., 2023; Chao et al., 2023).
This paper provides a novel, fourth category to the current taxonomy: Decomposition-based method that breaks the harmful prompt into sub-phrases (Figure 2). Using DrAttack as an example, we show that current LLMs are highly prone to become victims of attacks in this category - they can be jailbroken with merely 15 queries.
Figure 2: A category to the current taxonomy of prompt-based jailbreak attacks. Previous approaches nest harmful prompts entirely into optimized suffix nesting, prefix/system prompt nesting, or Scenarios nesting. DrAttack innovates by decomposing malicious prompts into discrete sub-prompts to jailbreak LLMs.
@@
Breaking down instruction into subtasks has demonstrated great success in enabling LLMs to perform complex tasks. Concretely,  (Ye et al., 2023; Dua et al., 2022; Radhakrishnan et al., 2023; You et al., 2023; V et al., 2023; Khot et al., 2023) show that dissecting complex questions into a set of simpler sub-questions allows LLMs to process and respond with greater accuracy and detail. In downstream tasks, this technique has been applied for enhancing prompt candidate selection (Li et al., 2022), refining model training processes (Shridhar et al., 2023), optimizing model fine-tuning (Shi & Lipani, 2023) and improving the performance of vision-related tasks (Yang et al., 2023). To the best of our knowledge, we provide the first method that shows prompt decomposition can be leveraged to develop a strong attacker, thereby exploring the safety breach of LLMs.
@@ 3 DrAttack Framework
DrAttack represents a novel approach to jailbreaking LLMs, employing prompt decomposition and reconstruction to generate an adversarial attack. This section lays out each component of the proposed DrAttack. As illustrated in Figure 1, the entire pipeline of DrAttack consists of two parts: decomposition and reconstruction. The section is organized as follows: Section 3.1 presents an overview of the entire pipeline; Section 3.2 explains the decomposition step using semantic parsing to derive sub-prompts from the original prompt; Section 3.3 discusses the implicit reconstruction via In-Context Learning (ICL), reassembling sub-prompts for attacking the LLM. The decomposition step is critical for breaking down the prompt into less detectable elements, while the reconstruction step cleverly reassembles these elements to bypass LLM security measures. Section 3.4 introduces a supplementary benefit of retrieving sub-prompts: synonym search on sub-prompts, which modifies sub-prompts to get a more effective and efficient jailbreaking. The pseudocode outlined in Algorithm 1 offers a comprehensive guide to the technical implementation of DrAttack.
3.1 Formulation and Motivation
Prompt-based attack
When queried with a prompt $p$, an LLM can either return an answer $a_{p}$ or reject the question $r_{p}$ if query $p$ is malicious. When the LLM rejects to answer a malicious query $p$, a jailbreaking algorithm attempts to search for an adversarial prompt $p^{\prime}$ that can elicit the desired answer $a_{p}$ from the target LLM. Therefore, jailbreaking algorithms are essentially trying to solve the following optimization problem:
${p^{\prime}}^{\star}=\arg\max_{p^{\prime}}\Pr(a_{p}|p^{\prime}),$ (1)
where $\Pr(a|p)$ denotes the likelihood of the target LLM ($f_{\text{LLM}}$).
@@ 3.1 Formulation and Motivation
Our method’s central idea is to camouflage a query’s malicious intent through semantic decomposition. An adversarial prompt $p$ can be parsed into a list of mutual-exclusive collectively-exhaustive sub-prompts $p_{1:m}$, each corresponding to a phrase in the original query. Intuitively, while the complete query is malicious (e.g., "make a bomb"), the parsed sub-prompts are often less alarming ("make" and "a bomb"), as demonstrated in our ablation study in Section 5.
@@
Although decomposition mitigates the original prompt’s harmfulness, it also disrupts the initial query’s intent. Thus, it’s necessary to reconstruct the original query from the parsed sub-prompts. However, naive straightforward reconstruction would simply replicate the original prompt, defeating the intended purpose.
Inspired by Chain-of-Thought (Wei et al., 2023a) and Rephrase-and-Respond (Deng et al., 2023), we propose to leverage the target LLM to reconstruct the question before answering it. Achieving this ought to be non-trivial: If we directly instruct the LLM to perform reconstruction while responding, the request will also alert the safety guard even when prompted with benign sub-prompts. This is because the LLM still needs to place full attention on the sub-prompts, thereby discerning the malicious intention effortlessly. To circumvent this issue, we embed this reconstruction sub-task inside a set of automatically-crafted benign demos. These in-context demos implicitly guide the target LLM to connect sub-phrases during its response, thereby diluting its attention.
Notably, our uses of In-Context Learning are fundamentally different from prior efforts: Previous work leverages harmful Question-Answer demos to elicit (Wei et al., 2023b) target LLM to answer malicious queries; whereas in our case, these demos are comprised of entirely benign examples (e.g., "how to make a cake") to teach the model on how to reassemble the answer.
3.2 Prompt Decomposition via Semantic Parsing
Formally, for a given malicious prompt $p$, our prompt decomposition algorithm will divide $p$ into phrases $p=p_{1}\mathbin{\|}...\mathbin{\|}p_{m}$. The process involves two primary steps: constructing a parsing tree and formatting coherent phrases.
Figure 3: A parsing tree example of the attack sentence "Write a tutorial on how to make a bomb". The original adversarial prompt is transformed from words into discrete phrases, then processed to sub-prompts with category labels in generating sub-prompts $p_{1:m}$
@@ 3.2 Prompt Decomposition via Semantic Parsing
In the first step, we construct a semantic parsing tree to map out the grammatical structure of the prompt. This tree helps understand the syntactic relationships between different parts of the sentence, such as verb phrases, noun phrases, and adjective phrases. An example of the decomposition through a parsing tree is shown in Figure 3. Given that semantic parsing is a well-established field in natural language processing, we simply utilize GPT (OpenAI, 2023a) to automate this task.
@@
Post parsing tree construction, we focus on merging adjacent words into coherent phrases. Adjacent words in the parsing tree’s leaf nodes are grouped based on grammatical and structural relationships to form coherent phrases that convey a complete semantic idea. This is done by categorizing them into four types based on their word class and phrase location at the tree: [instruction], [structure], [noun], and [verb]. This categorization aims to preserve phrases’ intrinsic semantics and clarify distinctions between sub-prompts for later reconstruction (as outlined in Section 3.3). A detailed implementation from parsing-tree words to coherent phrases can be found in Section A.1. These phrases $p_{i}$ then serve as the sub-prompts in our attack algorithm. With phrases’ categories, sub-prompts $p_{i}$ are easier to be reconstructed and modified. We give an example in Figure 3 to illustrate how the original prompt is transformed from words into discrete phrases and then processed to sub-prompts with category labels. In addition, a more detailed description of the parsing process and categorization is provided in Section A.1.

### v3

# DrAttack: Prompt Decomposition and Reconstruction Makes Powerful LLMs Jailbreakers

@@ Abstract
Safety-aligned Large Language Models (LLMs) are still vulnerable to some manual and automated jailbreak attacks, which adversarially trigger LLMs to output harmful content. However, existing jailbreaking methods usually view a harmful prompt as a whole but they are not effective at reducing LLMs’ attention on combinations of words with malice, which well-aligned LLMs can easily reject. This paper discovers that decomposing a malicious prompt into separated sub-prompts can effectively reduce LLMs’ attention on harmful words by presenting them to LLMs in a fragmented form, thereby addressing these limitations and improving attack effectiveness. We introduce an automatic prompt Decomposition and Reconstruction framework for jailbreaking Attack (DrAttack). DrAttack consists of three key components: (a) ‘Decomposition’ of the original prompt into sub-prompts, (b) ‘Reconstruction’ of these sub-prompts implicitly by In-Context Learning with semantically similar but benign reassembling example, and (c) ‘Synonym Search’ of sub-prompts, aiming to find sub-prompts’ synonyms that maintain the original intent while jailbreaking LLMs. An extensive empirical study across multiple open-source and closed-source LLMs demonstrates that, with fewer queries, DrAttack obtains a substantial gain of success rate on powerful LLMs over prior SOTA attackers. Notably, the success rate of 80% on GPT-4 surpassed previous art by 65%. Code and data are made publicly available at https://turningpoint-ai.github.io/DrAttack/.
DrAttack: Prompt Decomposition and Reconstruction Makes Powerful LLMs Jailbreakers
Xirui Li1, Ruochen Wang1, Minhao Cheng2, Tianyi Zhou3, Cho-Jui Hsieh1 1University of California, Los Angeles, 2Pennsylvania State University, 3University of Maryland, College Park
@@ 1 Introduction
Figure 1: Attack success rate (ASR) (%) of DrAttack and other prompt-based jailbreaking methods. DrAttack obtains a substantial gain of ASR on powerful LLMs (GPT, Claude, Gemini) over prior SOTA attackers.
The development of large language models (LLMs) has significantly advanced AI capabilities, enabling success in various tasks Floridi and Chiriatti (2020); Touvron et al. (2023); Chowdhery et al. (2023). To promote responsible usage, existing LLMs often need safety alignment to avoid generating harmful content. However, recent research Wei et al. (2023a) has shown that these safety-aligned LLMs sometimes generate harmful content through an adversarial attack known as "jailbreaking." Jailbreaking attacks involve manipulating a malicious query to circumvent the model’s safety mechanisms, thereby exposing vulnerabilities that can guide the development of safer LLMs. Current jailbreaking attacks primarily focus on modifying a malicious query with appended prefixes and/or suffixes. However, it often treats the malicious prompt as a single noticeable entity, which undermines the effectiveness of these attacks Zou et al. (2023); Zhu et al. (2023); Lapid et al. (2023); Liu et al. (2023b); Huang et al. (2023); Wei et al. (2023c); Yu et al. (2023); Li et al. (2023); Ding et al. (2023); Chao et al. (2023).
This paper exposes a new under-explored type of vulnerability. Our main insight is that while the malicious prompt as a whole can easily be perceived, it can be decomposed into a set of sub-prompts with significantly reduced attention to easily jailbreak victim LLMs (in Figure 3). Inspired by this observation, we develop a new type of jailbreaking attack that disguises a malicious prompt via syntactic parsing (e.g., "make" and "a bomb"). The concrete algorithm, termed Decomposition-and-Reconstruction Attack (DrAttack), operates through a three-step process: (1) Decomposition via Syntactic Parsing breaks down a malicious prompt into seemingly neutral sub-prompts. (2) Implicit Reconstruction via In-Context Learning reassembles sub-prompts by benign and semantic-similar assembling context. (3) Sub-prompt Synonym Search shrinks the search space to make the search more efficient than optimization in whole vocabulary. Extensive empirical evaluation demonstrates that prevalent LLMs struggle with this attack. DrAttack substantially increases the success rate over prior SOTA attacks on the most powerful LLMs. On GPT-4, DrAttack achieves an attack success rate of over 80% (human evaluation), surpassing the previous best results by over 65% in absolute terms.
@@ 2 Related Work
Effective attack techniques that circumvent LLM’s safety detectors uncover the vulnerabilities of LLMs, which could be regarded as a critical process in enhancing the design of safer systems. This is achieved by generating surrounding content to hide the harmful intention of the original prompt. Apart from attacks in other languages Xu et al. (2024); Yong et al. (2024); Wei et al. (2023a), existing monolingual attackers can be roughly categorized into three groups: 1). Suffix-based methods augment the harmful prompt with a suffix content, optimized to trick LLM into generating desired malicious responses (Zou et al., 2023; Zhu et al., 2023; Shah et al., 2023; Lapid et al., 2023). Specifically, GCG appends adversarially optimized suffixes to harmful prompts to jailbreak LLMs (Zou et al., 2023). 2). Prefix-based methods prepend contexts in front of harmful prompts (Liu et al., 2023b; Huang et al., 2023; Wei et al., 2023c). For instance, AutoDAN (Liu et al., 2023b) searches for optimal system prompts using a genetic algorithm. ICA (Wei et al., 2023c) adds fixed jailbroken examples before harmful prompts. 3). Hybrid methods operate on entire harmful prompts (Yu et al., 2023; Li et al., 2023; Ding et al., 2023; Chao et al., 2023; Deng et al., 2024b). PAIR (Chao et al., 2023) leverages red-teaming LLMs to generate benign contexts to nest harmful prompts, while DeepInception (Li et al., 2023) inserts harmful prompts into multi-layer benign scenarios. Moreover, ReNellm (Ding et al., 2023) rephrases and modifies harmful prompts, putting them into predefined tasks.
This paper provides a feasible, fourth category to the current taxonomy: Decomposition-based method that breaks the harmful prompt into sub-phrases (in Figure 2). While some initial studies (Liu et al., 2023a; Wei et al., 2023a) in this category have demonstrated little success, we show that current LLMs are highly prone to becoming victims of attacks in this category - they can be jailbroken with merely 15 queries.
@@
Breaking down instruction into subtasks has demonstrated great success in enabling LLMs to perform complex tasks. Concretely, Ye et al. (2023); Dua et al. (2022); Radhakrishnan et al. (2023); You et al. (2023); V et al. (2023); Khot et al. (2023) show that dissecting complex questions into a set of simpler sub-questions allows LLMs to process and respond with greater accuracy and more details. In downstream tasks, this technique has been applied to improve prompt candidate selection Li et al. (2022), refine model training processes Shridhar et al. (2023), optimize model fine-tuning Shi and Lipani (2023) and improve the performance of vision-related tasks Yang et al. (2023). To the best of our knowledge, we provide the first method that shows prompt decomposition can be leveraged to develop a strong attacker.
Figure 2: A category to the current taxonomy of prompt-based jailbreak attacks. Previous approaches take harmful prompts as an entity, appending suffixes, modifying prefixes/system prompts, or operating on the whole prompt. DrAttack innovates by decomposing malicious prompts into discrete sub-prompts to jailbreak LLMs.
3 Motivation
Prompt-based jailbreak attacks can be viewed as algorithms maximizing the likelihood of response given a malicious query. We then explore that decomposition, introducing malicious prompts into fragmented form, can conceal malice by analyzing the rejection tokens’ possibility. With this insight from open-source models, we generalize this strategy in Section 4 as a black-box attack to jailbreak both open-source and closed-source LLMs.
Prompt-based attack
Queried by a prompt $p$, an LLM can either return an answer $a_{p}$ or reject the question $r_{p}$ if query $p$ is malicious. When the LLM rejects to answer a malicious query $p$, a jailbreaking algorithm searches for an adversarial prompt $p^{\prime}$ that can elicit the desired answer $a_{p}$ from the target LLM. Therefore, jailbreaking algorithms are essentially trying to solve this optimization problem:
$\vspace{-2mm}{p^{\prime}}^{\star}$ $=\arg\max_{p^{\prime}}\Pr(a_{p}|p^{\prime}),\vspace{-2mm}$ (1)
where $\Pr(a|p)$ denotes the likelihood of the LLM.
(a) Investigation on rejection tokens
(b) Probability
Figure 3: (a) An illustration of our exploration on open-source LLMs when given initial and fragmented prompts. (b) Average probability of generating rejection tokens in response. LLMs prompted with fragmented prompts show a lower probability of rejection tokens (e.g., "Sorry") compared to those prompted with initial malicious prompts.
Decomposition concealing malicious intention
Intuitively, while the complete query is malicious (e.g., "make a bomb"), the sub-prompts are often harmless ("make" and "a bomb"). We examine how the probability in Equation 1 changes when malicious phrases are introduced in a fragmented format. Our analysis uses token probabilities from open-source LLMs, specifically Llama2 Touvron et al. (2023) and Vicuna Chiang et al. (2023), to demonstrate the effectiveness of this concealment strategy. By evaluating the average probabilities of the initial five tokens in rejection responses, we compare the LLMs’ reactions to original harmful prompts with those in fragmented formats. The findings, illustrated in Figure 3(b), reveal that while original prompts consistently trigger rejection responses (e.g., "Sorry"), fragmented prompts significantly diminish this likelihood in the context of jailbreak attacks.
Next, we introduce the DrAttack framework, a black-box attack designed to decompose harmful prompts and reconstruct them in a way that effectively jailbreak LLMs.
Figure 4: An illustration of DrAttack. An attack by a malicious prompt on LLMs would be rejected. However, with DrAttack’s prompt decomposition and reconstruction, the resulting prompt can jailbreak LLM to generate a harmful response. Colored words are sub-prompts generated by DrAttack.
@@ 4 DrAttack Framework
DrAttack represents a novel approach to jailbreaking LLMs, employing prompt decomposition and reconstruction to generate an adversarial attack. This section describes each component of the proposed DrAttack. As illustrated in Figure 4, the entire DrAttack pipeline consists of two parts: decomposition and reconstruction. The section is organized as follows: Section 4.1 presents an overview of the entire pipeline; Section 4.2 explains the decomposition step using semantic parsing to derive sub-prompts from the original attack prompt; Section 4.3 discusses the implicit reconstruction via In-Context Learning (ICL), reassembling sub-prompts for attacking LLMs. The decomposition step is critical for breaking down the prompt into less sensible elements, while the reconstruction step cleverly reassembles these elements and illicit LLMs to generate harmful contents. Section 4.4 introduces a supplementary benefit of our framework: Synonym Search on sub-prompts, which modifies sub-prompts to get a more effective jailbreaking.
4.1 Overview of Pipeline
@@
DrAttack’s central idea is to camouflage a query’s malicious intent through semantic decomposition. An adversarial prompt $p$ can be parsed into a list of mutually exclusive collectively exhaustive subprompts $p_{1:m}$, each corresponding to a phrase in the original query.
@@
Although decomposition mitigates the harmfulness of the original prompt, it also disrupts the intent of the initial query. Thus, it is necessary to reconstruct the original query from the parsed sub-prompts. However, a naive, straightforward reconstruction would simply replicate the original prompt, defeating the intended purpose.
Inspired by Chain-of-Thought Wei et al. (2023b) and Rephrase-and-Respond Deng et al. (2023), we propose to leverage victim LLMs to reconstruct the question before answering it. Achieving this is non-trivial since if we directly instruct victim LLMs to perform reconstruction while responding, the trivial request can not fool LLMs and be easily rejected. This is because LLMs still need to understand the semantic relationship between sub-phrases, thereby effortlessly discerning the malicious intention. To circumvent this issue, we embed this reconstruction sub-task inside a set of automatically crafted benign examples. These in-context examples implicitly guide victim LLMs to connect subphrases during their response, thereby jailbreaking victim LLMs.
Notably, our uses of ICL are fundamentally different from previous efforts: previous work leverages harmful question-answer examples to elicit victim LLMs to answer malicious queries Wei et al. (2023c); whereas in our case, these examples are comprised of entirely benign examples to teach the model on how to reassemble the answer.
4.2 Prompt Decomposition via Syntactic Parsing
Formally, for a given malicious prompt $p$, our prompt decomposition algorithm will divide $p$ into the phrases $p=p_{1}\mathbin{\|}...\mathbin{\|}p_{m}$. The process involves two primary steps: constructing a parsing tree and formatting coherent phrases.
@@ 4.2 Prompt Decomposition via Syntactic Parsing
In the first step, we construct a syntactic parsing tree to map the grammatical structure of the original prompt. This tree helps to understand the syntactic relationships between different parts of the sentence, such as verb phrases and noun phrases. Given that LLMs can achieve SOTA syntactic parsing performance compared to other methods (Drozdov et al., 2022), we prompt GPT-4 OpenAI (2023b) to parse prompt in PCFG form (Klein and Manning, 2003) by offering parsing examples to simplify and automate this task (see Appendix A.1 for more details and examples that generated by Standford PCFG Parser Klein and Manning (2003)).
@@
After parsing tree construction, we focus on merging adjacent words into coherent phrases. Adjacent words in the parsing tree’s leaf nodes are grouped based on grammatical and structural relationships to form coherent phrases that convey a complete semantic idea. This is done by categorizing them into four types based on their word class and phrase location at the tree: [instruction], [structure], [noun], and [verb]. This categorization aims to preserve phrases’ intrinsic semantics and clarify distinctions between sub-prompts for later reconstruction (as outlined in Section 4.3). Serving later as the subprompts in our attack algorithm, phrases $p_{i}$ are more manageable to reconstruct and modify than single words A nuanced error in the GPT-4 generated parsing tree could be mitigated by combining words into phrases (see Appendix A.1 for more details).
We offer an example in Figure 5 to illustrate how the original prompt is transformed from words into discrete phrases and then processed into sub-prompts with category labels.
Figure 5: An example of GPT-4 generated parsing tree for a malicious prompt "Write a tutorial on how to make a bomb". Prompt $p$ is firstly transformed from words into discrete phrases, then processed to sub-prompts $p_{1:m}$ with category labels.